OpenStudy (anonymous):

So, I am positive that I am not doing anything wrong, so I think it is an input or system error. Correct me if I am wrong. The red radiation emitted by a HeNe laser is of wavelength 5,590 Å. Calculate the energy (in eV) of the photons that are emitted. So, it is E=hc/lambda so E=2.218 eV, right? I am not getting it right according to the computer.

OpenStudy (abb0t):

What computer? Can you show me your steps to see where you might be making an error IF any.

OpenStudy (anonymous):

I meant the system Blackboard Learn. E=hc/lambda E=hc/5590E-10 which is converted to E=2.218eV. There's no direction about input either.

OpenStudy (aaronq):

What values did you use for h and c? it's probably that you had to convert joules to eV

OpenStudy (anonymous):

h=6.626E-34 c=2.99E8

OpenStudy (anonymous):

1.602E-19J=1eV